OpenStudy (anonymous):

find the taylor series for f centered at 4 if...

OpenStudy (anonymous):

\[f ^{n}(4) = (-1)^{n}n!/3^{n}(n+1)\]

OpenStudy (zarkon):

\[\sum_{n=0}^{\infty}\frac{f^{(n)}(4)(x-4)^n}{n!}\] replace \(f^{(n)}(4)\)

OpenStudy (anonymous):

replace it with what?

OpenStudy (lalaly):

with the expression you wrote above

OpenStudy (anonymous):

okay but how did he get that equation?

OpenStudy (lalaly):

The definition of a taylor series is

OpenStudy (anonymous):

okay how did he get (x-4)^n

OpenStudy (anonymous):

ohhhh nvm got it

OpenStudy (lalaly):

j\[\sum_{n=0}^{\infty}\frac{f^{(n)}(a)}{n!}(x-a)^n\]

OpenStudy (lalaly):

lol ok

OpenStudy (anonymous):

so the answer should be: \[\sum_{0}^{\infty} (-1)^{n}(x-4)^{n}\] ?

OpenStudy (anonymous):

actually no i messed up

OpenStudy (anonymous):

\[\sum_{0}^{\infty} (-1)^{n}(x-4)^{n}/3 ^{n}(n+1)\]

OpenStudy (amistre64):

assuming: \[f ^{n}(4) = (-1)^{n}\frac{n!}{3^{n}(n+1)}\] and given the generic from Zarkon: \[\sum_{n=0}^{\infty}\frac{f^{(n)}(4)}{n!}(x-4)^n\] lets combine the two of them \[\sum_{n=0}^{\infty}\ (-1)^{n}\frac{\frac{n!}{3^{n}(n+1)}}{n!}(x-4)^n\] \[\sum_{n=0}^{\infty}\ (-1)^{n}\frac{1}{3^{n}(n+1)}(x-4)^n\]

OpenStudy (amistre64):

looks good to me